OpenStudy (yash2651995):

f(x)=x^(1/x) f'(x) = ? and how?

OpenStudy (shubhamsrg):

well take natural log on both sides we have ln(fx) = (ln x)/x now differentiate both sides : (f'x)/(fx) = -(ln x)/x^2 + 1/x^2 we have fx =x^(1/x) just solve for f'x .. was i clear enough ?

OpenStudy (yash2651995):

thanks btw :)

OpenStudy (shubhamsrg):

glad to help ^_^

OpenStudy (yash2651995):

is it [1-lnx] x^((1/x) -2) ?

OpenStudy (shubhamsrg):

didnt get what you're asking ?

OpenStudy (shubhamsrg):

ohh..you're asking the final ans?

OpenStudy (shubhamsrg):

am really sorry,,too lazy to do that.. :P but if do it well enough,,should be correct :)

OpenStudy (yash2651995):

i need to find global maxima of fx that's why i was asking it.. =D

OpenStudy (shubhamsrg):

just put your f'x =0 and to make sure if its maxima,,check for f''x (gets bit complicated though)

OpenStudy (yash2651995):

ya i got it its maxima is at x=e try making its graph.. it was fun. thanks =)

OpenStudy (shubhamsrg):

glad to help ^_^